### v3

@@ Abstract
In recent decades, several assistive technologies have been developed to improve the ability of blind and visually impaired individuals to navigate independently and safely. At the same time, simultaneous localization and mapping (SLAM) techniques have become sufficiently robust and efficient to be adopted in developing these assistive technologies. We present the first systematic literature review of 54 recent studies on SLAM-based solutions for blind and visually impaired people, focusing on literature published from 2017 onward. This review explores various localization and mapping techniques employed in this context. We discuss the advantages and limitations of these techniques for blind and visually impaired navigation. Moreover, we examine the major challenges described across studies. We explain how SLAM technology offers the potential to improve the ability of visually impaired individuals to navigate effectively. Finally, we present future opportunities and challenges in this domain.
@@ I-B Contribution
This paper presents a systematic literature review (SLR) addressing fundamental questions regarding SLAM-based approaches for BVI navigation. This review provides insights into technological diversity, advantages, limitations, and the potential to address real-world challenges. While recognizing the broad range of potential research questions, such as sensors employed in studies, evaluation techniques, and human–computer interaction for feedback, we narrowed our focus to the four questions outlined in Table II. The primary contributions of this study are as follows:
  • Identification of SLAM Approaches: We systematically identify and categorize the diverse SLAM approaches adopted in the development of assistive systems tailored for visually impaired navigation.
@@
  • Classification of challenges: We identify and categorize studies that address challenging conditions relevant to SLAM-based navigation systems for the visually impaired.
  • Exploration of the potential for enhancing BVI navigation: We show that SLAM-based solutions have the potential to enhance the mobility and navigation capabilities of individuals with visual impairment.
@@ II SLR methodology
Element
Description [53]
Value
Population
The problem or situation
Visually impaired navigation
Intervention
The technology, tool or method under study
SLAM
Comparison (optional)
The technology, tool or method with which the intervention is compared
Outcome
Results that Intervention could produce
Lightweight, affordable, accurate, efficient assistive technology
Context
The specific context of the study
Autonomous mobility of visually impaired people
@@ II-B Planning the review
The first step in conducting SLR is to establish a protocol. The protocol outlines the review procedures and ensures replicability. Within the protocol, we formulated our research questions, designed a search strategy, and defined the specific criteria for selecting relevant studies. In addition, we defined a set of criteria presented in Table VI to evaluate the quality of the selected literature. Furthermore, to facilitate the extraction of data in alignment with our research questions, we designed a data-extraction form.
@@ II-B1 Research Questions and Motivation
No.
Research question
Description
RQ1
What localization and mapping approaches are used for the navigation of the visually impaired?
The target is to identify different localization and mapping techniques adopted for the development of assistive devices for visually impaired navigation.
RQ2
What are the advantages and limitations of SLAM techniques for BVI navigation?
The objective is to summarize the advantages and constraints of SLAM-based approaches for visually impaired navigation.
RQ3
What challenging situations have been addressed?
The purpose of this question is to know which of the challenging conditions (e.g. crowded environment, changing view point, challenging light conditions, dynamic objects, etc.) relevant to navigation systems have been considered.
RQ4
How does the proposed solution improve navigation using SLAM for individuals with impaired vision?
The research question seeks to understand how SLAM techniques can enhance mobility and navigation in individuals with visual impairments.
@@ II-B2 Search strategy
A key step in performing SLR is to design an effective search strategy. This strategy should be executed with reasonable effort to retrieve relevant studies from digital libraries [54]. The exhaustive search process in systematic reviews is a critical factor distinguishing them from traditional literature reviews [54], leading to a wider and more precise understanding of the topic under review. To design the search string presented in Figure 2 we first extracted keywords from the PICOC elements, including population, intervention, and outcomes. We then determined synonyms for each keyword to broaden the search string. The list of keywords and their synonyms related to each PICOC element is listed in Table III.
Figure 2: Search string
@@
Keywords
Synonyms
PICOC element
Visually impaired navigation
Blind navigation, Navigation assistance for the visually impaired, Navigational guidance for individuals with visual impairments, Navigational support for the visually disabled, Orientation and mobility for the visually impaired, Sight-impaired navigation, Visual impairment navigation aid, low vision navigation, partially sighted navigation
Population
SLAM
Real-time mapping and positioning, Simultaneous Localization and Mapping, Simultaneous mapping and position tracking, mapping and localization
Intervention
Accurate, efficient, reliable assistive technology
high-performance, precise, effective, trustworthy assistive technology
Outcome
@@
Digital source
Web address
# of papers
Last access date
ACM Digital Library
http://portal.acm.org
4
22 Jul 2023
Google Scholar
https://scholar.google.com
0
23 Jul 2023
IEEE Xplore
http://ieeexplore.ieee.org
11
22 Jul 2023
MDPI
https://www.mdpi.com
2
23 Jul 2023
PubMed
https://www.ncbi.nlm.nih.gov/pubmed
3
22 Jul 2023
Science Direct
http://www.sciencedirect.com
2
22 Jul 2023
Scopus
http://www.scopus.com
22
23 Jul 2023
Springer Link
http://link.springer.com
2
22 Jul 2023
Taley & Francis
https://www.tandfonline.com
0
22 Jul 2023
Wiley Online Library
https://onlinelibrary.wiley.com
1
22 Jul 2023
We utilized the Advanced Search feature in digital libraries to gain more control over our search parameters. The title, abstract, and keyword fields were selected to retrieve the search results. Searching on Google Scholar is somewhat different from searching for other digital libraries. Unlike other platforms, Google Scholar does not suggest various filters, requiring the manual incorporation of filters into the search string. Additionally, to identify English-written studies, we adjusted the language preference settings within our Google Scholar account to filter the search results in English. Upon receiving the ”Use fewer boolean connectors (max 8 per field)” message on Sciencedirect, we divided search string into separate queries to keep the number of boolean connectors within the limit. We then executed these queries separately, and merged the results to obtain a set of relevant studies.
@@ II-B3 Selection criteria
Criteria
Inclusion
Exclusion
Availability
Available in full text
Not accessible in specific databases
Language
English
Not written in English
Publication period
From 2017 to July 2023
Prior 2017
Type of source
Conferences and journals papers
Books, dissertation, newsletters, speeches, technical reports, white papers
Relevance
Papers relevant to at least two research questions
Outside the scope of our research
@@ II-C1 Identification
During the initial phase of our review, we conducted searches across the digital libraries using custom-formulated queries for each library. The initial searches yielded a large number of papers irrelevant to our research questions, resulting in 6,809 records. As explained in Section II-B2, we employed a single query in all libraries, except ScienceDirect and Google Scholar, to retrieve relevant records. The results obtained from these searches were exported in the BibTex format, a process facilitated by the export citation features available in the libraries. The BibTex data were then imported into the Parsifal framework for subsequent stages of our review. Springer and Google Scholar do not provide direct options for exporting data in BibTex format. To address this issue, we used Zotero and its browser plugin, Zotero Connector, to streamline the process. With these tools, we added paper information from webpage views to Zotero and subsequently retrieved BibTex data.
@@ II-C2 Study selection
In the next step, we performed a fast reading of the full text of the remaining papers, excluding 265 studies that were outside the scope of our research. We then evaluated the quality of the studies based on the quality assessment criteria defined in the SLR protocol. Five studies were removed during the assessment of study quality. Table VI lists the quality assessment criteria for our SLR.
TABLE VI: Quality assessment criteria and weights
Criteria
Weight
Is there an adequate description of the context in which the research was carried out?
0.0, 0.5, 1.0
Does the methodology take into account both localization and mapping issues?
0.0, 0.1, 2.0
Is the solution proposed well presented?
0.0, 0.5, 1.0
Is there a clear statement of findings?
0.0, 0.5, 1.0
Is the research design appropriate to address the aims of the research?
0.0, 0.5, 1.0
Does the study add value to the research community?
0.0, 0.5, 1.0
@@
[…]
@@
It is important to note that the last search conducted in digital libraries was on July 23, 2023, and that for forward snowballing was on August 12, 2023. These dates should be considered as the starting points for future reviews. The publications included in our review are listed in Tables VIIIX and categorized based on their publication venues. These tables provide an overview of the literature, including the paper title, author names, publication year, location, and source through which the studies were discovered. Among the 54 studies included in our SLR, 27 were sourced from journals, as presented in Table VII. The remaining 27 studies were presented at conferences, as shown in Tables VIII and IX.
@@ II-C3 Data extraction
  • Short summary of the paper: A concise overview of the main points and findings of the study
@@
TABLE VII: List of journal publications included in SLR
@@
[56]
Sonification of navigation instructions for people with visual impairment
@@
Dragan Ahmetovic and Federico Avanzini and Adriano Baratè and Cristian Bernareggi and Marco Ciardullo and Gabriele Galimberti and Luca A. Ludovico and Sergio Mascetti and Giorgio Presti
International Journal of Human-Computer Studies
2023
Science@Direct
[57]
@@
Sensing and Navigation of Wearable Assistance Cognitive Systems for the Visually Impaired
Li, Guoxin and Xu, Jiaqi and Li, Zhijun and Chen, Chao and Kan, Zhen
IEEE Transactions on Cognitive and Developmental Systems
2023
IEEE Digital Library
[58]
Mixture reality-based assistive system for visually impaired people
@@
Jucheng Song and Jixu Wang and Shuliang Zhu and Haidong Hu and Mingliang Zhai and Jiucheng Xie and Hao Gao
Displays
2023
Science@Direct
[59]
@@
Research on Design and Motion Control of a Considerate Guide Mobile Robot for Visually Impaired People
Zhang, Bin and Okutsu, Mikiya and Ochiai, Rin and Tayama, Megumi and Lim, Hun-Ok
IEEE Access
2023
Scopus
[60]
UNav: An Infrastructure-Independent Vision-Based Navigation System for People with Blindness and Low Vision
@@
Yang, Anbang and Beheshti, Mahya and Hudson, Todd E and Vedanthan, Rajesh and Riewpaiboon, Wachara and Mongkolwat, Pattanasak and Feng, Chen and Rizzo, John-Ross
Sensors
2022
PubMed
[61]
Multi-Floor Indoor Localization Based on Multi-Modal Sensors
@@
Zhou, Guangbing and Xu, Shugong and Zhang, Shunqing and Wang, Yu and Xiang, Chenlu
Sensors
2022
Scopus
[62]
@@
Knowledge driven indoor object-goal navigation aid for visually impaired people
Hou, Xuan and Zhao, Huailin and Wang, Chenxu and Liu, Huaping
Cognitive Computation and Systems
2022
Wiley Online Library
[63]
@@
Indoor-Guided Navigation for People Who Are Blind: Crowdsourcing for Route Mapping and Assistance
Plikynas, Darius and Indriulionis, Audrius and Laukaitis, Algirdas and Sakalauskas, Leonidas
Applied Sciences (Switzerland)
2022
Scopus
[64]
@@
A Multi-Sensory Guidance System for the Visually Impaired Using YOLO and ORB-SLAM
Xie, Zaipeng and Li, Zhaobin and Zhang, Yida and Zhang, Jianan and Liu, Fangming and Chen, Wei
Information
2022
MDPI
[65]
@@ A Multi-Sensory Guidance System for the Visually Impaired Using YOLO and ORB-SLAM
Egocentric Human Trajectory Forecasting with a Wearable Camera and Multi-Modal Fusion
Qiu, Jianing and Chen, Lipeng and Gu, Xiao and Lo, Frank P.-W. and Tsai, Ya-Yen and Sun, Jiankai and Liu, Jiaqi and Lo, Benny
IEEE Robotics and Automation Letters
2022
Scopus
[66]
@@
A wearable navigation device for visually impaired people based on the real-time semantic visual slam system
Chen, Zhuo and Liu, Xiaoming and Kojima, Masaru and Huang, Qiang and Arai, Tatsuo
Sensors
2021
PubMed
[67]
@@
Multimodal sensing and intuitive steering assistance improve navigation and mobility for people with impaired vision
Slade, Patrick and Tambe, Arjun and Kochenderfer, Mykel J.
Science Robotics
2021
Scopus
[68]
@@
Assistive Navigation Using Deep Reinforcement Learning Guiding Robot With UWB/Voice Beacons and Semantic Feedbacks for Blind and Visually Impaired People
Lu, Chen-Lung and Liu, Zi-Yan and Huang, Jui-Te and Huang, Ching-I and Wang, Bo-Hui and Chen, Yi and Wu, Nien-Hsin and Wang, Hsueh-Cheng and Giarré, Laura and Kuo, Pei-Yi
Frontiers in Robotics and AI
2021
Scopus
[69]
@@
Indoor Wearable Navigation System Using 2D SLAM Based on RGB-D Camera for Visually Impaired People
Hakim, Heba and Fadhil, Ali
Advances in Intelligent Systems and Computing
2021
Scopus
[70]
@@
An RGB-D Camera Based Visual Positioning System for Assistive Navigation by a Robotic Navigation Aid
Zhang, He and Jin, Lingqiu and Ye, Cang
IEEE/CAA Journal of Automatica Sinica
2021
IEEE Digital Library
[71]
Hierarchical visual localization for visually impaired people using multimodal images
@@
Cheng, Ruiqi and Hu, Weijian and Chen, Hao and Fang, Yicheng and Wang, Kaiwei and Xu, Zhijie and Bai, Jian
Expert Systems with Applications
2021
Scopus
[72]
@@
Indoor Topological Localization Based on a Novel Deep Learning Technique
Liu, Qiang and Li, Ruihao and Hu, Huosheng and Gu, Dongbing
Cognitive Computation
2020
Scopus
[73]
@@
Combining Obstacle Avoidance and Visual Simultaneous Localization and Mapping for Indoor Navigation
Jin, SongGuo and Ahmed, Minhaz Uddin and Kim, Jin Woo and Kim, Yeong Hyeon and Rhee, Phill Kyu
Symmetry
2020
MDPI
[74]
@@
Wearable travel aid for environment perception and navigation of visually impaired people
Bai, Jinqiang and Liu, Zhaoxiang and Lin, Yimin and Li, Ye and Lian, Shiguo and Liu, Dijun
Electronics (Switzerland)
2019
Scopus
[75]
@@
An ARCore based user centric assistive navigation system for visually impaired people
Zhang, Xiaochen and Yao, Xiaoyu and Zhu, Yi and Hu, Fei
Applied Sciences (Switzerland)
2019
Scopus
[76]
@@
Virtual-Blind-Road Following-Based Wearable Navigation Device for Blind People
Bai, Jinqiang and Lian, Shiguo and Liu, Zhaoxiang and Wang, Kai and Liu, Dijun
IEEE Transactions on Consumer Electronics
2018
IEEE Digital Library
[77]

### v4

@@ Abstract
In recent decades, several assistive technologies have been developed to improve the ability of blind and visually impaired individuals to navigate independently and safely. At the same time, simultaneous localization and mapping (SLAM) techniques have become sufficiently robust and efficient to be adopted in developing these assistive technologies. We present the first systematic literature review of 54 recent studies on SLAM-based solutions for blind and visually impaired people, focusing on literature published from 2017 onward. This review explores various localization and mapping techniques employed in this context. We systematically identified and categorized diverse SLAM approaches and analyzed their localization and mapping techniques, sensor types, computing resources, and machine-learning methods. We discuss the advantages and limitations of these techniques for blind and visually impaired navigation. Moreover, we examine the major challenges described across studies, including practical considerations that affect usability and adoption. Our analysis also evaluates the effectiveness of these SLAM-based solutions in real-world scenarios and user satisfaction, providing insights into their practical impact on BVI mobility. The insights derived from this review identify critical gaps and opportunities for future research activities, particularly in addressing the challenges presented by dynamic and complex environments. We explain how SLAM technology offers the potential to improve the ability of visually impaired individuals to navigate effectively. Finally, we present future opportunities and challenges in this domain.
@@ I Introduction
This SLR is designed to act as a resource for the academic and research communities. The objective of this review is to explore and highlight the strengths and potential limitations of the current SLAM applications for visually impaired navigation. This study aims to inform and guide subsequent research. The insights derived from this review identify critical gaps and opportunities for future research, particularly for tackling the challenges presented by dynamic and complex environments. Such environments pose unique difficulties for visually impaired navigation, and addressing them through advanced SLAM technologies could lead to significant improvements in both the effectiveness and reliability of assistive solutions.
@@ I-B Contribution
This paper presents a systematic literature review (SLR) addressing fundamental questions regarding SLAM-based approaches for BVI navigation. This review provides insights into technological diversity, advantages, limitations, and the potential to address real-world challenges. While recognizing the broad range of potential research questions we narrowed our focus to the four questions outlined in Table II. The primary contributions of this study are as follows:
  • Identification of SLAM approaches: We systematically identified and categorized the diverse SLAM approaches adopted in the development of assistive systems tailored for visually impaired navigation. This includes analyzing the localization and mapping techniques, sensor types, computing resources, and machine-learning methods used in these approaches.
@@
  • Classification of challenges: We identify and categorize studies that address challenging conditions relevant to SLAM-based navigation systems for the visually impaired. In addition, we discuss practical considerations that affect the usability and adoption of these systems.
  • Exploration of the potential for enhancing BVI navigation: We analyzed how the proposed SLAM-based approaches improved navigation in visually impaired individuals. In addition, we evaluated the effectiveness of these solutions in real-world scenarios and assessed user satisfaction to understand their practical impact on BVI mobility.
@@ II SLR methodology
Element Description [53] Value
Population The problem or situation Visually impaired navigation
Intervention The technology, tool or method under study SLAM
Comparison (optional) The technology, tool or method with which the intervention is compared -
Outcome Results that Intervention could produce Lightweight, affordable, accurate, efficient assistive technology
Context The specific context of the study Autonomous mobility of visually impaired people
@@ II-B Planning the review
The first step in conducting SLR is to establish a protocol. The protocol outlines the review procedures and ensures replicability. Within the protocol, we formulated our research questions, designed a search strategy, and defined the specific criteria for selecting relevant studies. In addition, we defined a set of criteria presented in Table VIII to evaluate the quality of the selected literature. Furthermore, to facilitate the extraction of data in alignment with our research questions, we designed a data-extraction form.
@@ II-B1 Research Questions and Motivation
No. Research question Description
RQ1 What localization and mapping approaches are used for the navigation of the visually impaired? The target is to identify different localization and mapping techniques adopted for the development of assistive devices for visually impaired navigation.
RQ2 What are the advantages and limitations of SLAM techniques for BVI navigation? The objective is to summarize the advantages and constraints of SLAM-based approaches for visually impaired navigation.
RQ3 What challenging situations have been addressed? The purpose of this question is to know which of the challenging conditions (e.g. crowded environment, changing view point, challenging light conditions, dynamic objects, etc.) relevant to navigation systems have been considered.
RQ4 How does the proposed solution improve navigation using SLAM for individuals with impaired vision? The research question seeks to understand how SLAM techniques can enhance mobility and navigation in individuals with visual impairments.
@@ II-B2 Search strategy
A key step in performing SLR is to design an effective search strategy. This strategy should be executed with reasonable effort to retrieve relevant studies from digital libraries [54]. The exhaustive search process in systematic reviews is a critical factor distinguishing them from traditional literature reviews [54], leading to a wider and more precise understanding of the topic under review. To design the search string we first extracted keywords from the PICOC elements, including population, intervention, and outcomes. We then determined synonyms for each keyword to broaden the search string. The list of keywords and their synonyms related to each PICOC element is listed in Table III.
@@
Keywords Synonyms PICOC element
Visually impaired navigation Blind navigation, Navigation assistance for the visually impaired, Navigational guidance for individuals with visual impairments, Navigational support for the visually disabled, Orientation and mobility for the visually impaired, Sight-impaired navigation, Visual impairment navigation aid, low vision navigation, partially sighted navigation Population
SLAM Real-time mapping and positioning, Simultaneous Localization and Mapping, Simultaneous mapping and position tracking, mapping and localization Intervention
Accurate, efficient, reliable assistive technology high-performance, precise, effective, trustworthy assistive technology Outcome
@@
Digital source Web address # of papers Last access date
ACM Digital Library http://portal.acm.org 4 22 Jul 2023
Google Scholar https://scholar.google.com 0 23 Jul 2023
IEEE Xplore http://ieeexplore.ieee.org 11 22 Jul 2023
MDPI https://www.mdpi.com 2 23 Jul 2023
PubMed https://www.ncbi.nlm.nih.gov/pubmed 3 22 Jul 2023
Science Direct http://www.sciencedirect.com 2 22 Jul 2023
Scopus http://www.scopus.com 22 23 Jul 2023
Springer Link http://link.springer.com 2 22 Jul 2023
Taley & Francis https://www.tandfonline.com 0 22 Jul 2023
Wiley Online Library https://onlinelibrary.wiley.com 1 22 Jul 2023
We utilized the Advanced Search feature in digital libraries to gain more control over our search parameters. The title, abstract, and keyword fields were selected to retrieve the search results. Searching on Google Scholar is somewhat different from searching for other digital libraries. Unlike other platforms, Google Scholar does not suggest various filters, requiring the manual incorporation of filters into the search string. Additionally, to identify English-written studies, we adjusted the language preference settings within our Google Scholar account to filter the search results in English.
@@ II-B3 Selection criteria
Criteria Inclusion Exclusion
Availability Available in full text Not accessible in specific databases
Language English Not written in English
Publication period From 2017 to July 2023 Prior 2017
Type of source Conferences and journals papers Books, dissertation, newsletters, speeches, technical reports, white papers
Relevance Papers relevant to at least two research questions Outside the scope of our research
@@ II-C1 Identification
During the initial phase of our review, we conducted searches across the digital libraries using custom-formulated queries for each library. For each dataset, we ran three different search strings (SS1, SS2, and SS3), as shown in Table VI, consisting of various combinations of keywords, booleans, and wildcard operators. These search strings were applied to all digital libraries except Google Scholar. For Google Scholar, we initially used keywords similar to those used in SS1, resulting in over 11,000 results. Upon reviewing a subset of these, we determined that a significant number were not relevant to our topic. Consequently, we decided to use only the primary keywords (shown in Table III) to construct the search string for this digital library.
We selected the search string that yielded the most results to identify primary studies and then applied exclusion criteria to the results obtained from the search strings used for each library. We observed that SS3, which incorporated ’Orientation and mobility’ to refine the search by focusing on more specialized literature, did not yield better results than SS1, which included the general term ’mobility’, across all digital libraries. This indicates that the broader term ’mobility’ was sufficient to capture the necessary literature. The specificity of SS3 did not contribute to additional relevant results. Additionally, upon receiving the message ’Use fewer Boolean connectors (maximum 8 per field)’ while running SS1 on ScienceDirect, we switched to SS2 to maintain the number of Boolean connectors within the limit.
The initial searches of all digital libraries resulted in 6,809 records. The search strings used for each digital library is presented in Table VII.
TABLE VI: Search strings applied to digital libraries, featuring keywords and operators to identify primary studies.
SS1 SS2 SS3 SS4
(”Visually impaired” OR blind OR ”visual impairment*” OR ”visually disabled” OR ”Sight impaired”) AND (navigation* OR mobility) AND (SLAM OR ”Simultaneous Localization and Mapping” OR ”Real-time mapping and positioning” OR ”Simultaneous mapping and position tracking” OR ”mapping and localization”) AND (technology OR aid OR support OR assist*) (”Visually impaired” OR blind) AND (navigation OR mobility) AND (SLAM OR ”Simultaneous Localization and Mapping”) (”Visually impaired” OR blind OR ”visual impairment*” OR ”visually disabled” OR ”Sight Impaired”) AND (navigation* OR ”Orientation and mobility”) AND (SLAM OR ”Simultaneous Localization and Mapping” OR ”Real-time mapping and positioning” OR ”Simultaneous mapping and position tracking” OR ”mapping and localization”) AND (technology OR aid OR support OR assist*) (”conference paper” OR ”journal”) AND (”Visually impaired”) AND (navigation) AND (SLAM) AND (-review) AND ( -survey)
TABLE VII: Utilization of search strings by digital libraries
Digital source Search string Number of results
ACM Digital Library SS1 284
Google Scholar SS4 602
IEEE Xplore SS2 50
MDPI SS2 191
PubMed SS1 8
Science Direct SS2 518
Scopus SS2 1510
Springer Link SS2 2585
Taley & Francis SS2 486
Wiley Online Library SS2 575
The results obtained from digital libraries searches were exported in the BibTex format, a process facilitated by the export citation features available in the libraries. The BibTex data were then imported into the Parsifal framework for subsequent stages of our review. Springer and Google Scholar do not provide direct options for exporting data in BibTex format. To address this issue, we used Zotero and its browser plugin, Zotero Connector, to streamline the process. With these tools, we added paper information from webpage views to Zotero and subsequently retrieved BibTex data.
@@ II-C2 Study selection
In the next step, we performed a fast reading of the full text of the remaining papers, excluding 265 studies that were outside the scope of our research. We then evaluated the quality of the studies based on the quality assessment criteria defined in the SLR protocol. Five studies were removed during the assessment of study quality. Table VIII lists the quality assessment criteria for our SLR.
TABLE VIII: Quality assessment criteria and weights
Criteria Weight
Is there an adequate description of the context in which the research was carried out? 0.0, 0.5, 1.0
Does the methodology take into account both localization and mapping issues? 0.0, 0.1, 2.0
Is the solution proposed well presented? 0.0, 0.5, 1.0
Is there a clear statement of findings? 0.0, 0.5, 1.0
Is the research design appropriate to address the aims of the research? 0.0, 0.5, 1.0
Does the study add value to the research community? 0.0, 0.5, 1.0
@@
Sensitivity=Number of relevant studies retrievedTotal number of relevant $\text{Sensitivity}=\dfrac{\text{Number of relevant studies retrieved}}{\text{Total number of relevant studies}}\times 100$
@@
It is important to note that the last search conducted in digital libraries was on July 23, 2023, and that for forward snowballing was on August 12, 2023. These dates should be considered as the starting points for future reviews. The publications included in our review are listed in Tables IXXI and categorized based on their publication venues. These tables provide an overview of the literature, including the paper title, author names, publication year, location, and source through which the studies were discovered. Among the 54 studies included in our SLR, 27 were sourced from journals, as presented in Table IX. The remaining 27 studies were presented at conferences, as shown in Tables X and XI.
Additionally, Tables XIIXV summarize the perspectives and innovations presented in the publications with insights into their limitations and advantages. These tables demonstrate the research issues addressed and the contributions of each study, highlighting the strengths and potential drawbacks of the proposed solutions. They also indicate which solutions are open source, with only seven papers having some or all parts of the project available as open source. Links to the sources are provided in these tables if they are directly available in the relevant papers.
@@ II-C3 Data extraction
  • Short summary of the paper: A concise overview of the main points and findings of the study.
  • Research issue and contribution: Summary of the research issues addressed and contributions of the studies.
@@
  • Localization and mapping accuracy and robustness: Assessment of accuracy and robustness levels in localization and mapping techniques.
  • Running time: Analysis of the running time for localization and mapping techniques.
@@
  • Computing resources: Identification of computing resources used in SLAM-based solutions.
@@
  • Practical challenges and operational efficiency: Evaluation of the user-friendliness, cost-efficiency, weight, comfort for extended use, adjustable fit, fatigue mitigation, and portability of the SLAM-based assistive tools.
  • System prototype information: Detailed information on functionalities, sensors, computing resources, human-computer interaction (HCI) mechanisms, assistive tools, and battery life.
  • User evaluation: Assessment of user satisfaction of the SLAM-based assistive tools.
  • Machine learning techniques: Identification of machine learning techniques used in assistive solutions.
  • Open-source availability: Identification of open-source contributions in the reviewed studies.
@@
TABLE IX: List of journal publications included in SLR
@@
[56] Sonification of navigation instructions for people with visual impairment Dragan Ahmetovic and Federico Avanzini and Adriano Baratè and Cristian Bernareggi and Marco Ciardullo and Gabriele Galimberti and Luca A. Ludovico and Sergio Mascetti and Giorgio Presti International Journal of Human-Computer Studies 2023 Science@Direct
@@
[57] Sensing and Navigation of Wearable Assistance Cognitive Systems for the Visually Impaired Li, Guoxin and Xu, Jiaqi and Li, Zhijun and Chen, Chao and Kan, Zhen IEEE Transactions on Cognitive and Developmental Systems 2023 IEEE Digital Library
@@
[58] Mixture reality-based assistive system for visually impaired people Jucheng Song and Jixu Wang and Shuliang Zhu and Haidong Hu and Mingliang Zhai and Jiucheng Xie and Hao Gao Displays 2023 Science@Direct
@@
[59] Research on Design and Motion Control of a Considerate Guide Mobile Robot for Visually Impaired People Zhang, Bin and Okutsu, Mikiya and Ochiai, Rin and Tayama, Megumi and Lim, Hun-Ok IEEE Access 2023 Scopus
@@
[60] UNav: An Infrastructure-Independent Vision-Based Navigation System for People with Blindness and Low Vision Yang, Anbang and Beheshti, Mahya and Hudson, Todd E and Vedanthan, Rajesh and Riewpaiboon, Wachara and Mongkolwat, Pattanasak and Feng, Chen and Rizzo, John-Ross Sensors 2022 PubMed
@@
[61] Multi-Floor Indoor Localization Based on Multi-Modal Sensors Zhou, Guangbing and Xu, Shugong and Zhang, Shunqing and Wang, Yu and Xiang, Chenlu Sensors 2022 Scopus
@@
[62] Knowledge driven indoor object-goal navigation aid for visually impaired people Hou, Xuan and Zhao, Huailin and Wang, Chenxu and Liu, Huaping Cognitive Computation and Systems 2022 Wiley Online Library
@@
[63] Indoor-Guided Navigation for People Who Are Blind: Crowdsourcing for Route Mapping and Assistance Plikynas, Darius and Indriulionis, Audrius and Laukaitis, Algirdas and Sakalauskas, Leonidas Applied Sciences (Switzerland) 2022 Scopus
@@
[64] A Multi-Sensory Guidance System for the Visually Impaired Using YOLO and ORB-SLAM Xie, Zaipeng and Li, Zhaobin and Zhang, Yida and Zhang, Jianan and Liu, Fangming and Chen, Wei Information 2022 MDPI
@@
[65] Egocentric Human Trajectory Forecasting with a Wearable Camera and Multi-Modal Fusion Qiu, Jianing and Chen, Lipeng and Gu, Xiao and Lo, Frank P.-W. and Tsai, Ya-Yen and Sun, Jiankai and Liu, Jiaqi and Lo, Benny IEEE Robotics and Automation Letters 2022 Scopus
@@
[66] A wearable navigation device for visually impaired people based on the real-time semantic visual slam system Chen, Zhuo and Liu, Xiaoming and Kojima, Masaru and Huang, Qiang and Arai, Tatsuo Sensors 2021 PubMed
@@
[67] Multimodal sensing and intuitive steering assistance improve navigation and mobility for people with impaired vision Slade, Patrick and Tambe, Arjun and Kochenderfer, Mykel J. Science Robotics 2021 Scopus
@@
[68] Assistive Navigation Using Deep Reinforcement Learning Guiding Robot With UWB/Voice Beacons and Semantic Feedbacks for Blind and Visually Impaired People Lu, Chen-Lung and Liu, Zi-Yan and Huang, Jui-Te and Huang, Ching-I and Wang, Bo-Hui and Chen, Yi and Wu, Nien-Hsin and Wang, Hsueh-Cheng and Giarré, Laura and Kuo, Pei-Yi Frontiers in Robotics and AI 2021 Scopus
@@
[69] Indoor Wearable Navigation System Using 2D SLAM Based on RGB-D Camera for Visually Impaired People Hakim, Heba and Fadhil, Ali Advances in Intelligent Systems and Computing 2021 Scopus
@@
[70] An RGB-D Camera Based Visual Positioning System for Assistive Navigation by a Robotic Navigation Aid Zhang, He and Jin, Lingqiu and Ye, Cang IEEE/CAA Journal of Automatica Sinica 2021 IEEE Digital Library
@@
[71] Hierarchical visual localization for visually impaired people using multimodal images Cheng, Ruiqi and Hu, Weijian and Chen, Hao and Fang, Yicheng and Wang, Kaiwei and Xu, Zhijie and Bai, Jian Expert Systems with Applications 2021 Scopus
@@
[72] Indoor Topological Localization Based on a Novel Deep Learning Technique Liu, Qiang and Li, Ruihao and Hu, Huosheng and Gu, Dongbing Cognitive Computation 2020 Scopus
@@
[73] Combining Obstacle Avoidance and Visual Simultaneous Localization and Mapping for Indoor Navigation Jin, SongGuo and Ahmed, Minhaz Uddin and Kim, Jin Woo and Kim, Yeong Hyeon and Rhee, Phill Kyu Symmetry 2020 MDPI
@@
[74] Wearable travel aid for environment perception and navigation of visually impaired people Bai, Jinqiang and Liu, Zhaoxiang and Lin, Yimin and Li, Ye and Lian, Shiguo and Liu, Dijun Electronics (Switzerland) 2019 Scopus
@@
[75] An ARCore based user centric assistive navigation system for visually impaired people Zhang, Xiaochen and Yao, Xiaoyu and Zhu, Yi and Hu, Fei Applied Sciences (Switzerland) 2019 Scopus
@@
[76] Virtual-Blind-Road Following-Based Wearable Navigation Device for Blind People Bai, Jinqiang and Lian, Shiguo and Liu, Zhaoxiang and Wang, Kai and Liu, Dijun IEEE Transactions on Consumer Electronics 2018 IEEE Digital Library